2 The standard reaction Gibbs energy for K:CrO4(aq) + 2Ag(s) + 2FeCl3(aq) → Ag;CrO4(s) + 2FeCl2(aq) + 2KCI(aq) is-62.5 kJ mol at 298 K. Calculate the standard emfof the corresponding galvanic cell and (a) (b) the standard potential of the Ag:CrO:] Ag, Cro couple.

media%2F6e6%2F6e6c5a8f-ca41-4ef4-927b-55


Please can I get step solution to question. Answer is provided.
Answer #1

K2CrO4 + 2Ag + 2FeCl3 à Ag2CrO4 + 2FeCl2 +2KCl

2Ag + CrO42- → Ag2CrO4 + 2e-

2FeCl3 + 2e- → 2FeCl2 + 2Cl-

delta G = -nFEcell

-62.5 x 103 J/mol= -(2 mol) (96500 c/mol) Ecell

Ecell = 62.5/(2 x 96.5) = 0.324 V

Ecell = Ecello – RT/nF lnq

0.324 = RT/nF ln [Red/Oxid]

0.324 = SPEFe -SPEAg

Fe+3/Fe2+ = 0.77

SPEAg = 0.77-0.324 = 0.45V